A boat moves through the water with two forces acting on it. One is a 1,800-N forward push by the water on the propeller, and the other is a 1,200-N resistive force due to the water around the bow.

(a) What is the acceleration of the 1,300-kg boat?
m/s 2

(b) If it starts from rest, how far will the boat move in 22.5 s?
m

(c) What will its velocity be at the end of that time?
m/s
Answer #1

(a) Acceleration a = Fnet/m

= (1800 - 1200)/1300

= 0.462 m/s2

(b) S = 1/2 * at2

= 1/2*0.462*22.52

= 116.83 m

(c) v = at

= 0.462*22.5

= 10.4 m/s

Kindly upvote :)